# A2PO: Towards Effective Offline Reinforcement Learning from an Advantage-aware Perspective

### v2

@@ 2 Related Works
Advantage-weighted Offline RL Method employs weighted sampling to prioritize training transitions with high advantage values from the offline dataset. To enhance sample efficiency, Peng et al. [28] introduce an advantage-weighted maximum likelihood loss by directly calculating advantage values via trajectory return. [26] further use the critic network to estimate advantage values for advantage-weighted policy training. This technique has been incorporated as a subroutine in other works [18, 46] for agent policy extraction. Recently, AW methods have also been well studied in addressing the constraint conflict issue that arises from the mixed-quality dataset [4, 52, 36]. Several studies present advantage-weighted behavior cloning as a direct objective function [52] or an explicit policy constraint [29]. [4] propose the Latent Advantage-Weighted Policy Optimization (LAPO) framework, which employs an advantage-weighted loss to train CVAE for generating high-advantage actions based on the state condition. Besides AW methods, Hong et al. [15] enhance the classical offline RL training objective with the weight of subsequent return. On the other hand, Hong et al. [16] directly learning the optimal policy density as the weight function to enable sampling from high-performing policies. However, this AW mechanism inevitably diminishes the data diversity in the offline dataset. In contrast, our A2PO directly conditions the agent policy on both the state and the estimated advantage value, enabling effective utilization of all samples with varying quality.
@@ 4.2 Agent Policy Optimization
With this form, the advantage-aware policy $\pi_{\omega}$ is expected to produce action with different qualities that are positively correlated to the designated advantage input $\xi$ which is normalized within $(-1,1)$ in Eq. (4). Therefore, the output optimal action $a^{*}$ is obtained by $c^{*}=s\,||\,\xi^{*}$ input with $\xi^{*}=1$. It should be noted that the critic networks are to approximate the expected value of the optimal policy $\pi_{\omega}(\cdot|c^{*})$. The agent optimization, following the actor-critic framework, encompasses policy evaluation and policy improvement steps. During the policy evaluation step, the critic is optimized through the minimization of the temporal difference loss with the optimal policy $\pi_{\omega}(\cdot|c^{*})$, as follows:
$\mathcal{L}_{\text{TD}}(\theta,\phi)=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim\mathcal{D},\\ \tilde{z}^{*}\sim\pi_{\omega}(\cdot|c^{*}),\\ a^{*}_{\xi}\sim p_{\psi}(\cdot|\tilde{z}^{*},c^{*})\end{subarray}}\Big{[}\sum_{i}\big{[}r+\gamma V_{\hat{\phi}}(s)-Q_{\theta_{i}}(s,a)\big{]}^{2}+\big{[}r+\gamma\min_{i}Q_{\hat{\theta}_{i}}(s^{\prime},a^{*}_{\xi})-V_{\phi}(s)\big{]}^{2}\Big{]},$ (7)
where $Q_{\hat{\theta}}$ and $V_{\hat{\phi}}$ are the target networks updated softly. The first term of $\mathcal{L}_{\text{TD}}$ is to optimize the Q-network while the second term is for the V-network. Different from Eq. (1), we introduce the target Q and V simultaneously to stablily optimize the mutual online network.
@@ 5 Experiments
To illustrate the effectiveness of the proposed A2PO method, we conduct experiments on the D4RL benchmark [7]. We aim to answer the following questions: (1) Can A2PO outperform the advanced offline RL methods in both the single-quality datasets and mixed-quality datasets? (Section 5.2) (2) How do different components of A2PO contribute to the overall performance? (Section 5.3 and Appendix CF) (3) How does A2PO perform under mixed-quality datasets with varying single-quality samples? (Section 6 and Appendix G) (4) Can the A2PO agent effectively estimate the quality of different transitions? (Section 5.4 and Appendix H) (5) How does the time overhead of A2PO compare to other baselines? (Section 5.6)
@@ 5.1 Experiment Settings
Comparison Methods and Hyperparameters. We compare the proposed A2PO to several advanced offline RL methods: BCQ [10], TD3+BC [8], CQL [20], EQL [46], especially the advantage-weighted offline RL methods: AWAC [26], IQL [18], CQL+AW [15], LAPO [4]. Besides, we also select the vanilla BC method [30], the model-based offline RL method MOPO [49], and the emerging diffusion-based method Diffusion-QL [43], for comparison. We report the performance of baselines using the best results reported from their own paper. The detailed hyperparameters of A2PO are given in Appendix B.2.
@@
medium expert halfcheetah 55.2 89.1 90.7 91.6 63.3 90.6 94.3 42.8 86.7 89.4 94.2 95.6 $\pm$
@@
antmaze-umaze-diverse 45.6 55.0 71.4 84.0 0.0 65.4 24.0 49.3 62.2 36.0 84.1 93.3 $\pm$
4.7
antmaze-medium-diverse 0.0 0.0 3.0 53.7 0.0 70.0 46.4 0.7 70.0 6.0 1.2 86.7 $\pm$
9.4
antmaze-large-diverse 0.0 2.2 0.0 14.9 0.0 42.5 56.6 1.0 47.5 2.0 0.0  53.3
@@
4.7
Maze Total  47.9 147.5 260.6 175.8 3.1 327.7 410.6 220.8 307.3 96.5 276.2 637.9
@@ 5.2 Comparison on D4RL Benchmarks
Figure 3: Test return difference of A2PO with different discrete advantage conditions during training compared with original A2PO with continuous advantage condition during training. Task abbreviations are listed in Appendix B.1. Test returns are reported in Appendix C.
@@ 5.3 Ablation Analysis
Figure 4: Learning curves of A2PO under different fixed advantage inputs during the test while using the original continuous advantage condition for training. Test returns are reported in Appendix D.
@@ 5.4 Visualization
Figure 5 presents the visualization of A2PO latent representation. The uniformly sampled advantage condition $\xi$ combined with the initial state $s$, are fed into the actor-network to get the latent representation generated by the final layer of the actor. The result demonstrates that the representations converge according to the advantage and the actual return. Moreover, upon comparing Figure 5(a,b), as well as Figure 5(c,d) separately, we observe that the latent action representation follows the same alteration pattern based on the actual real return. These observations demonstrate that our advantage-aware policy effectively capture policies with different returns by the designated advantage input $\xi$. This provides compelling evidence for the efficacy of the A2PO policy construction. More experiments of advantage estimation conducted on different tasks and datasets are presented in Appendix H.
@@ 5.5 Robustness
Figure 6: Compare the returns of A2PO under random-expert dataset with different high-quality data proportions $\sigma$ in the Gym tasks. Detail returns are reported in Appendix G.
@@ Appendix A Pseudocode
         Optimize CVAE encoder $q_{\varphi}$ and decoder $p_{\psi}$ according to Eq. (5) as
@@
     Optimize critic networks $Q_{\theta}$ and $V_{\phi}$ according to Eq. (7) as
$\mathcal{L}_{\text{TD}}(\theta,\phi)=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim\mathcal{D},\\ \tilde{z}^{*}\sim\pi_{\omega}(\cdot|c^{*}),\\ a^{*}_{\xi}\sim p_{\psi}(\cdot|\tilde{z}^{*},c^{*})\end{subarray}}\bigg{[}\sum_{i}\big{[}r+\gamma V_{\hat{\phi}}(s)-Q_{\theta_{i}}(s,a)\big{]}^{2}+\big{[}r+\gamma\min_{i}Q_{\hat{\theta}_{i}}(s^{\prime},a^{*}_{\xi})-V_{\phi}(s)\big{]}^{2}\bigg{]}.$
     Optimize actor network $\pi_{\omega}$ according to Eq. (8) as
@@ B.1 Task Abbreviations and Task Versions
In order to improve the readability and conciseness, we adopt abbreviations for the Gym tasks throughout the main text. The corresponding abbreviations for each task are provided in Table 3. For the specific task versions, we use the ‘-v2’ version for the Gym tasks, the ‘-v1’ version for the Maze2d tasks, the ‘-v2’ version for the Antmaze tasks, the ‘-v0’ version for the Kitchen tasks, the ‘-v1’ version for the Adroit tasks.
@@ B.2 Implementation Details
In this section, we provide the implementation details of our experiments. We conducted our experiments using PyTorch 3.8 [27] on a cluster of 8 A6000 GPUs. Each run required approximately 8 hours to complete 1 million steps. The source code will be made publicly available upon the publication of this paper.
The critic, actor, and CVAE networks are all constructed using a 2-layer MLP with 256 hidden units. They are updated using the Adam optimizer with a learning rate of $3\times 10^{-4}$. Additionally, we employ a target critic network with a soft-update rate of $5\times 10^{-3}$. Following the TD+3BC approach [8], we incorporate Q normalization, policy noise, and policy clipping during the training process. For the hyperparameter value $\alpha$ in Q normalization, following the strategy of Wang et al. [43], we set $\alpha=5.0$ for Antmaze tasks to emphasize strong and stable Q Learning. For other tasks, setting $\alpha=1.0$ produces state-of-the-art results. By performing a simple hyperparameter search on CVAE step $K$, we found that the optimal settings are $4\times 10^{5}$ for maze2d-umaze and maze2d-medium tasks, and $2\times 10^{5}$ is robust enough for the other tasks. Among the tasks, the continuous form of the advantage signal $\xi$, as described in Eq. (4), consistently yields excellent results. The only exception is the maze2d-umaze task, where utilizing a discrete form with $\epsilon=0.1$ achieves similarly high performance. Further details on the sensitivity of the $\xi$ type can be found in Section 5.3 and Appendix C.
@@
Appendix C Analysis of Advantage Condition Input for Training
@@ 90.6 $\pm$
Appendix D Analysis of Advantage Condition Input for Testing
In this section, we provide a full comparison of different fixed advantage condition inputs for testing on Gym tasks in Table 5 as a supplement for Figure 4. It should be noted that in contrast to the ablation study shown in Appendix C which investigates different methods for computing advantages during CVAE and agent learning, this section utilizes the original continuous advantage computation methods described in Eq. (4). However, the designated $\xi$ is varied during the testing process. On one hand, the CVAE decoder $p_{\psi}$ can generate CVAE-approximated behavior policies by varying the input $\xi$. On the other hand, Eq. (8) for policy improvement regulates the advantage-aware policy to closely resemble samples from the same advantage condition. This indicates that the quality of the generated action should be positively correlated with the specified action input. The experimental outcomes demonstrate the relationship between varying values of $\xi$ and the corresponding A2PO policy returns: the agent performance consistently improves as the fixed advantage condition input $\xi$ increases. Moreover, the gap between the returns increases as the offline dataset includes more diverse behavior policies. These observations serve as strong evidence for the effectiveness of A2PO in effectively disentangling the behavior policies within a multi-quality dataset.
@@ 97.7 $\pm$
Appendix E Analysis of the CVAE Policy
The CVAE policy corresponds to the CVAE decoder $p_{\psi}(a|z_{0},c^{*})$, where $z_{0}$ is sampled from $\mathcal{N}(0,1)$, state-advantage, $\xi^{*}=1$ represents the largest advantage condition, and $c^{*}=s\,||\,\xi^{*}$. As described in Section 4.1, the advantage-aware CVAE utilizes the advantage condition computed by the agent critic to construct the ELBO loss, which is formulated as advantage-guided supervised learning [12]. The CVAE decoder $p_{\psi}(a|z_{0},c^{*})$ generates action outputs of varying qualities based on the target advantage input signal $\xi$. By conditioning on the maximum normalized advantage value $\xi^{*}$ to get the optimal action of CVAE and agent policy, we present thorough comparison results of the CVAE policy and agent policy in Table 6. The performance of the CVAE policy indicates that it only demonstrates superior performance in a limited set of tasks and datasets, specifically the hopper-random and walker2d-random environments. The A2PO agent consistently outperforms the CVAE agent in the majority of cases. This performance gap is particularly significant in the newly constructed mix-quality datasets including random-expert and random-medium-expert highlighting the conflict issue. These findings suggest that A2PO achieves well-disentangled behavior policies and an optimal agent policy, surpassing the capabilities of CVAE-reconstructed behavior policies.
@@
Appendix F Analysis of A2PO Policy Optimization
In this section, we consider the effectiveness of the BC regularization term in the A2PO policy optimization. Previous methods [51, 4] construct the agent policy in the CVAE latent space using the deterministic policy gradient loss in policy optimization. The BC regularization term is not utilized in these methods because the latent action space inherently imposes a constraint on the action output. However, in our advantage-aware policy, we incorporate the BC term into the actor loss to guarantee that the actions chosen by the policy align with the advantage condition $\xi$ determined by the critic. To assess the impact of our A2PO architecture and the BC regularization, we conduct an ablation study by eliminating the regularization term. In this case, both the A2PO and the LAPO method with advantage-weighting employ the same policy optimization loss of deterministic policy gradient. The comparison results are shown in Table 7. The results indicate that even without the BC regularization term, A2PO consistently outperforms LAPO in the majority of tasks. Moreover, the BC term in A2PO can enhance its performance in most cases. This comparison highlights the superior performance achieved by A2PO, showcasing its effective disentanglement of action distributions from different behavior policies in order to enforce a reasonable advantage-aware policy constraint and obtain an optimal agent policy.
Table 7: Test returns of LAPO and our proposed A2PO without BC term in the policy optimization step. Bold indicates that the better performance among LAPO and A2PO w/o BC.
Source Task LAPO A2PO w/o BC  A2PO (Ours)

### v3

@@ 2 Related Works
Advantage-weighted Offline RL Method employs weighted sampling to prioritize training transitions with high advantage values from the offline dataset. To enhance sample efficiency, Peng et al. [28] introduce an advantage-weighted maximum likelihood loss by directly calculating advantage values via trajectory return. [26] further use the critic network to estimate advantage values for advantage-weighted policy training. This technique has been incorporated as a subroutine in other works [18, 47] for agent policy extraction. Recently, AW methods have also been well studied in addressing the constraint conflict issue that arises from the mixed-quality dataset [4, 54, 37]. Several studies present advantage-weighted behavior cloning as a direct objective function [54] or an explicit policy constraint [29]. [4] propose the Latent Advantage-Weighted Policy Optimization (LAPO) framework, which employs an advantage-weighted loss to train CVAE for generating high-advantage actions based on the state condition. Besides AW methods, Hong et al. [15] enhance the classical offline RL training objective with the weight of subsequent return. On the other hand, Hong et al. [16] directly learning the optimal policy density as the weight function to enable sampling from high-performing policies. However, this AW mechanism inevitably diminishes the data diversity in the dataset. In contrast, our A2PO directly conditions the agent policy on both the state and the estimated advantage value, enabling effective utilization of all samples with varying quality.
@@ 4.2 Agent Policy Optimization
With this form, the advantage-aware policy $\pi_{\omega}$ is expected to produce action with different qualities that are positively correlated to the designated advantage input $\xi$ which is normalized within $(-1,1)$ in Eq. (4). Therefore, the output optimal action $a^{*}$ is obtained by $c^{*}=s\,||\,\xi^{*}$ input with $\xi^{*}=1$. It should be noted that the critic networks are to approximate the expected value of the optimal policy $\pi_{\omega}(\cdot|c^{*})$. The agent optimization, following the actor-critic framework, encompasses policy evaluation and policy improvement steps. During the policy evaluation step, the critic is updated through the minimization of the temporal difference loss with the optimal policy $\pi_{\omega}(\cdot|c^{*})$. Specifically, for the V-network $v_{\phi}$, we employ the one-step Bellman operator to approximate the state value under the current agent-aware policy, conditioned on the optimal advantage input $\xi^{*}=1$, as follows:
$\mathcal{L}_{\text{TD}}(\phi)=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim\mathcal{D},\\ \tilde{z}^{*}\sim\pi_{\omega}(\cdot|c^{*}),\\ a^{*}_{\xi}\sim p_{\psi}(\cdot|\tilde{z}^{*},c^{*})\end{subarray}}\Big{[}r+\gamma\min_{i}Q_{\hat{\theta}_{i}}(s^{\prime},a^{*}_{\xi})-V_{\phi}(s)^{2}\Big{]},$ (7)
where $Q_{\hat{\theta}}$ is the target network updated softly. As for the Q-networks, both of the two Q-network entities $Q_{\theta_{i}}$ are optimized with agent policy $\pi_{\omega}(\cdot|c^{*})$ following Equation 1.
@@ 5 Experiments
To illustrate the effectiveness of the proposed A2PO method, we conduct experiments on the D4RL benchmark [7]. We aim to answer the following questions: (1) Can A2PO outperform the advanced offline RL methods in both the single-quality datasets and mixed-quality datasets? (Section 5.2) (2) How do different components of A2PO contribute to the overall performance? (Section 5.3 and Appendix DG) (3) How does A2PO perform under mixed-quality datasets with varying single-quality samples? (Section 6 and Appendix H) (4) Can the A2PO agent effectively estimate the quality of different transitions? (Section 5.4 and Appendix I) (5) How does the time overhead of A2PO compare to other baselines? (Section 5.6)
@@ 5.1 Experiment Settings
Comparison Methods and Hyperparameters. We compare the proposed A2PO to several advanced offline RL methods: BCQ [10], TD3+BC [8], CQL [20], EQL [47], especially the advantage-weighted offline RL methods: AWAC [26], IQL [18], CQL+AW [15], LAPO [4]. Besides, we also select the vanilla BC method [30], the model-based offline RL method MOPO [50], and the emerging diffusion-based method Diffusion-QL [44], for comparison. We report the performance of baselines using the best results reported from their own paper. More comparison results can be found in Appendix C. The detailed hyperparameters of A2PO are given in Appendix B.2.
@@
medium expert halfcheetah 55.2 89.1 90.7 91.6 63.3 90.6 96.8 42.8 86.7 84.0 94.2  95.6
$\pm$
@@
antmaze-umaze-diverse 45.6 55.0 71.4 84.0 0.0 50.8 66.2 49.3 62.2 54.0 0.0  72.6
@@
10.2
antmaze-medium-diverse 0.0 0.0 3.0 53.7 0.0 62.2 78.6 0.7 70.0 24.0 30.2 80.2 $\pm$
4.0
antmaze-large-diverse 0.0 2.2 0.0 14.9 0.0 38.0 56.6 1.0 47.5 40.0 22.3  52.1
$\pm$
7.9
Maze Total  47.9 147.5 260.6 175.8 3.1 300.8 485.0 220.8 307.3 170.5 243.4 609.5
@@ 5.2 Comparison on D4RL Benchmarks
Figure 3: Test return difference of A2PO with different discrete advantage conditions during training compared with original A2PO with continuous advantage condition during training. Task abbreviations are listed in Appendix B.1. Test returns are reported in Appendix D.
@@ 5.3 Ablation Analysis
Figure 4: Learning curves of A2PO under different fixed advantage inputs during the test while using the original continuous advantage condition for training. Test returns are reported in Appendix E.
@@ 5.4 Visualization
Figure 5 presents the visualization of A2PO latent representation. The uniformly sampled advantage condition $\xi$ combined with the initial state $s$, are fed into the actor-network to get the latent representation generated by the final layer of the actor. The result demonstrates that the representations converge according to the advantage and the actual return. Moreover, upon comparing Figure 5(a,b), as well as Figure 5(c,d) separately, we observe that the latent action representation follows the same alteration pattern based on the actual real return. These observations demonstrate that our advantage-aware policy effectively capture policies with different returns by the designated advantage input $\xi$. This provides compelling evidence for the efficacy of the A2PO policy construction. More experiments of advantage estimation conducted on different tasks and datasets are presented in Appendix I.
@@ 5.5 Robustness
Figure 6: Compare the returns of A2PO under random-expert dataset with different high-quality data proportions $\sigma$ in the Gym tasks. Detail returns are reported in Appendix H.
@@ Appendix A Pseudocode
         Optimize CVAE encoder $q_{\varphi}$ and decoder $p_{\psi}$ according to Eq. 5 as
@@
     Optimize critic networks $Q_{\theta}$ and $V_{\phi}$ according to Eq. 1 and Eq. 7 as
$\mathcal{L}_{\text{TD}}(\theta,\phi)=\mathbb{E}_{\begin{subarray}{c}(s,a,r,s^{\prime})\sim\mathcal{D},\\ \tilde{z}^{*}\sim\pi_{\omega}(\cdot|c^{*}),\\ a^{*}_{\xi}\sim p_{\psi}(\cdot|\tilde{z}^{*},c^{*})\end{subarray}}\bigg{[}\sum_{i}\big{[}r+\gamma\min_{j}Q_{\hat{\theta}_{j}}(s^{\prime},a^{*}_{\xi})-Q_{\theta_{i}}(s,a)\big{]}^{2}+\big{[}r+\gamma\min_{i}Q_{\hat{\theta}_{i}}(s^{\prime},a^{*}_{\xi})-V_{\phi}(s)\big{]}^{2}\bigg{]}.$
     Optimize actor network $\pi_{\omega}$ according to Eq. 8 as
@@ B.1 Task Abbreviations and Task Versions
In order to improve the readability and conciseness, we adopt abbreviations for the Gym tasks throughout the main text. The corresponding abbreviations for each task are provided in Table 3. For the specific task versions, we use the ‘-v2’ version for the Gym tasks, the ‘-v1’ version for the Maze2d tasks, the ‘-v0’ version for the Antmaze tasks, the ‘-v0’ version for the Kitchen tasks, the ‘-v1’ version for the Adroit tasks.
@@ B.2 Implementation Details
In this section, we provide the implementation details of our experiments. We conducted our experiments using PyTorch 3.8 [27] on a cluster of 4 A40 GPUs.The source code will be made publicly available upon the publication of this paper.
The critic, actor, and CVAE networks are all constructed using a 2-layer MLP with 256 hidden units. They are updated using the Adam optimizer with a learning rate of $3\times 10^{-4}$. Additionally, we employ a target critic network with a soft-update rate of $5\times 10^{-3}$. Following the TD+3BC approach [8], we incorporate Q normalization, policy noise, and policy clipping during the training process. or the hyperparameter value $\alpha$ in Q normalization, we adopted the strategy proposed by Wang et al. [44], selecting values of $5.0$ and $6.0$ for the antmaze-medium/large-diverse tasks respectively to promote robust and stable Q Learning. In contrast, we chose $\alpha=0.005$ for the kitchen-partial task to facilitate policy regularization. For other tasks, setting $\alpha=1.0$ yields state-of-the-art results. As for the CVAE step $K$, 0.2M steps demonstrate sufficient robustness for most tasks, while the optimal settings for the maze2d-umaze/medium tasks are 0.4M and 0.1M steps, respectively. Among the tasks, the continuous form of the advantage signal $\xi$, as described in Eq. (4), consistently produces outstanding results across most scenarios. However, the advantage distribution may vary significantly across all the tasks, necessitating the fine-tuning of advantage computation methods based on online evaluations. The maze2d-umaze task requires a discrete form with $\epsilon=0.1$. In the case of the maze2d-medium task, we utilize an $\epsilon$ value of $0.2$. For both the maze2d-large and antmaze-umaze tasks, we increase this to $\epsilon=0.8$. In contrast, for the kitchen-partial/mixed tasks, we adopt a value of $\epsilon=0.0$. Further details on this sensitivity can be found in Section 5.3 and Appendix D.
@@
Appendix C Additional Comparisons with More Baselines
In this section, we have added the data rebalance baseline ReD [51]; novel policy-constrain baseline PRDC [33] and return-conditioned baselines Decision Transformer and %BC [2] as the comparison baselines to further evaluate the superiority of our A2PO. The results are illustrated in Table 4. Our A2PO method demonstrates comparable performance to various offline RL methods. While the performance of A2PO is slightly inferior to that of PRDC in the medium and medium-replay dataset, it surpasses all additional baselines in the majority of scenarios.
Table 4: Test returns of our proposed A2PO and additional baselines on the locomotion tasks. The italic results of the baselines are obtained from the corresponding original paper.
Source Task %BC DT IQL+ReD PRDC A2PO (Ours)
medium halfcheetah 42.5 42.6 47.6 63.5  47.1
$\pm$
0.2
hopper 56.9 67.6 66.0 100.3  80.3
$\pm$
4.0
walker2d 75.0 74.0 78.6 85.2  84.9
$\pm$
0.2
medium replay halfcheetah 40.6 36.6 44.3 55.0  44.8
$\pm$
0.2
hopper 75.9 82.7 101.0 100.1 101.6 $\pm$
1.3
walker2d 62.5 66.6 79.5 92.0 82.8
$\pm$
1.7
medium expert halfcheetah 92.9 86.8 92.6 94.5 95.6 $\pm$
0.5
hopper 110.9 82.7 101.0 109.2 113.4 $\pm$
0.5
walker2d 109.0 108.1 110.5 111.2 112.1 $\pm$
0.2
random medium halfcheetah 40.1 42.0 42.0 56.5  48.5
$\pm$
0.3
hopper 21.0 3.1 6.7 5.5 62.1 $\pm$
2.8
walker2d 32.0 66.9 60.0 5.5 82.3 $\pm$
0.4
random expert halfcheetah 7.7 10.3 42.4 1.3 90.3 $\pm$
1.6
hopper 2.4 90.2 16.7 24.8 112.5 $\pm$
1.3
walker2d 53.4 103.4 93.2 1.1 109.1 $\pm$
1.4
random medium expert halfcheetah 29.1 42.6 39.1 10.5 90.6 $\pm$
1.6
hopper 62.0 46.1 31.3 88.5 107.8 $\pm$
0.4
walker2d 10.6 78.8 52.0 4.9 97.7 $\pm$
6.7
Appendix D Analysis of Advantage Condition Input for Training
@@ 90.6 $\pm$
Appendix E Analysis of Advantage Condition Input for Testing
In this section, we provide a full comparison of different fixed advantage condition inputs for testing on Gym tasks in Table 6 as a supplement for Figure 4. It should be noted that in contrast to the ablation study shown in Appendix D which investigates different methods for computing advantages during CVAE and agent learning, this section utilizes the original continuous advantage computation methods described in Eq. (4). However, the designated $\xi$ is varied during the testing process. On one hand, the CVAE decoder $p_{\psi}$ can generate CVAE-approximated behavior policies by varying the input $\xi$. On the other hand, Eq. (8) for policy improvement regulates the advantage-aware policy to closely resemble samples from the same advantage condition. This indicates that the quality of the generated action should be positively correlated with the specified action input. The experimental outcomes demonstrate the relationship between varying values of $\xi$ and the corresponding A2PO policy returns: the agent performance consistently improves as the fixed advantage condition input $\xi$ increases. Moreover, the gap between the returns increases as the offline dataset includes more diverse behavior policies. These observations serve as strong evidence for the effectiveness of A2PO in effectively disentangling the behavior policies within a multi-quality dataset.
@@ 97.7 $\pm$
Appendix F Analysis of the CVAE Policy
The CVAE policy corresponds to the CVAE decoder $p_{\psi}(a|z_{0},c^{*})$, where $z_{0}$ is sampled from $\mathcal{N}(0,1)$,and $c^{*}=s\,||\,\xi^{*}$. As described in Section 4.1, the advantage-aware CVAE utilizes the advantage condition computed by the agent critic to construct the ELBO loss, which is formulated as advantage-guided supervised learning [12]. The CVAE decoder $p_{\psi}(a|z_{0},c^{*})$ generates action outputs of varying qualities based on the target advantage input signal $\xi$. By conditioning on the maximum normalized advantage value $\xi^{*}$ to get the optimal action of CVAE and agent policy, we present thorough comparison results of the CVAE policy and agent policy in Table 7. The performance of the CVAE policy indicates that it only demonstrates superior performance in a limited set of tasks and datasets, specifically the hopper-random and walker2d-random environments. The A2PO agent consistently outperforms the CVAE agent in the majority of cases. This performance gap is particularly significant in the newly constructed mix-quality datasets including random-expert and random-medium-expert highlighting the conflict issue. These findings suggest that A2PO achieves well-disentangled behavior policies and an optimal agent policy, surpassing the capabilities of CVAE-reconstructed behavior policies.
@@
Appendix G Analysis of A2PO Policy Optimization
In this section, we evaluate the effectiveness of the policy regularization term within the policy improvement step as shown in Equation 8. Previous approaches [53, 4] that construct agent policies within the CVAE latent space do not incorporate the BC regularization term, as the latent action space inherently imposes constraints on the action outputs. In contrast, our advantage-aware policy integrates the BC term into the actor loss to explicitly ensure appropriate pessimism. Meanwhile, to guarantee the policy action outputs are consistent with the advantage condition $\xi$, we develop a variant of A2PO, designated A2PO$[a^{*}_{\text{cvae}}]$, which constrains the agent output to the optimal CVAE policy $p_{\psi}(\cdot|z_{0}\sim\mathcal{N}(0,I),c^{*})$. Unlike this variant, the original A2PO utilizes the MSE loss to align the agent action output $\pi(\cdot|\tilde{z}^{*},c^{*})$ with the sampled action $a$. To assess the impact of our A2PO architecture and the BC regularization, we conduct an ablation study by comparing the results of different A2PO variants and LAPO, as demonstrated in Table 8. The results indicate that even without the BC regularization term, A2PO consistently outperforms LAPO in most tasks. Furthermore, the original formulation of the BC term in the main text can enhance performance in most cases. This comparison highlights the superior performance achieved by A2PO, showcasing its effective disentanglement of action distributions from different behavior policies to enforce a reasonable advantage-aware policy constraint and obtain an optimal agent policy.
Table 8: Test returns of LAPO, A2PO and its variants in relation to policy regularization.
Source Task LAPO A2PO w/o BC A2PO$[a^{*}_{\text{cvae}}]$  A2PO (Ours)